3) A certain part of the electromagnetic spectrum ranges from 200 nm to 400 nm. What is the highest frequency associated with this portion of the spectrum? (c 3.00 x 108 m/s) A) 1.50 x 1014 Hz B) 7.50 x 1013 Hz E) 7.50 x 1015 Hz D) 1.50 x 1015 Hz 4) A radio station broadcasts at a frequency of 80 MHz. How far from the transmitter will this signal travel in 67 ms? (c 3.0x 108 m/s) A) 20 x 106 m B) 67 x 106 m C) 40 km D) 80 km
Answer #1

3) Option D) is correct

f = c/L = 3*10^8/(200*10^-9) = 1.5*10^15 Hz

4) Option A) is correct

No matter what the frequency speed = distance / time
so distance = speed * time

The speed of light is 3 * 10 ^ 8 and the time is 67 * 10 ^ -3

so distance= 200 * 10 ^ 5
= 2 * 10 ^7 m